I am trying to help my daughter with this question as well:
Malcom has a choice of profit sharing plans at work. He can receive 1/90 of the company's gross income or he can receive 1/60 of the company's net profit. Gross income is the total amount the company takes in. Profit is the income that is left after the expenses have been subtracted off. The company's expenses for one month are $100,000. What must the the gross income be for the two options to be equal? How much is each option worth at this point?

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

1/90x=1/60*100,000
x/90=100,000/60 now cross multiply
60x=90*100,000
x=9,000,000/60
x=$150,000 the amount of the gross to equal the profit offer.
proof
150,000/90=100,000/60
16.67=16.67

Answer by stanbon(48568)

Malcom has a choice of profit sharing plans at work. He can receive 1/90 of the company's gross income or he can receive 1/60 of the company's net profit. Gross income is the total amount the company takes in. Profit is the income that is left after the expenses have been subtracted off. The company's expenses for one month are $100,000. What must the the gross income be for the two options to be equal? How much is each option worth at this point?
----------------
Let gross sales be "x" dollars.
Net profit is "x-100,000"
---------------
Gross income plan: I(x) = (1/90)x
Net income plan: I(x) = (1/60)(x-100,000)
-------------
EQUATION:
Gross plan = Net plan
(1/90)x = (1/60)(x-100,000)
Multiply both sides by 180 to get:
2x = 3(x-100,000)
2x = 3x - 300,000
x = 300,000
Plans are even when gross sales is $300,000
=================
Cheers,
Stan H.